OpenStudy (anonymous):

Solve the equation for x. -7 – 10x = -9x

pooja195 (pooja195):

−7−10x=−9x −7+−10x=−9x −10x−7=−9x Add 9x to both sides. −10x−7+9x=−9x+9x −x−7=0 −x−7+7=0+7 −x=7 finish it off :)